OpenStudy (anonymous):

HELP! Find a line that goes through (1,-6) and is parallel to the line of x+2y=6

OpenStudy (0bucchr):

gradient = -0.5 y = mx + c -6 = (-0.5)*(1) + c c = -5.5 y = -0.5x - 5.5

OpenStudy (anonymous):

1) find the parallel slope you need: x+2y=6; so 2y=-x+6; so y=-(1/2)x+3 so the slope is -(1/2) 2) find a line that goes through (1,-6) with a slope of -(1/2) \[y=-\frac{1}{2}x+b\] \[-6=-\frac{1}{2}(1) + b\] \[-6+\frac{1}{2} =b\] Therefore the line is \[y=-\frac{1}{2}x -\frac{11}{2}\]

OpenStudy (anonymous):

in my textbook it says that the answer is x+2y+11=0. I just dont understand how to get that

OpenStudy (anonymous):

Thats the same thing. Multiply my whole equation by 2 you get \[2y=-x-11\] add x+11 to both sides gives x+2y+11-0

OpenStudy (anonymous):

Sorry gives x+2y+11=0

OpenStudy (anonymous):

oh wow. that was simple lol. thank you for the help :)